# KIX: A Knowledge and Interaction-Centric Metacognitive Framework for Task Generalization

### v3

@@ Abstract
People aptly exhibit general intelligence behaviors through flexible problem-solving and the ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time. In contrast, artificial agents tend to be specialists, lacking such generalist behaviors. To bridge this gap, artificial agents will require understanding and exploiting critical structured knowledge representations. We introduce a metacognitive reasoning framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects, by leveraging a type space, facilitate the learning of transferable interaction concepts and promote generalization. This framework offers a principled approach for integrating knowledge into reinforcement learning and holds promise as an enabler for generalist behaviors in artificial intelligence, robotics, and autonomous systems.
@@ Introduction
The ability to actively reason, flexibly evaluate options and learn from experience is a hallmark of human intelligence. Core to this intelligence is the human ability to identify natural decompositions within complex environments and tasks with ease, enabling flexible and generalist behaviors. Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue necessitate a knowledge and interaction-centric metacognitive approach.
Meta-reasoning about next actions, their values, and possible next states underpins flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition. This form of reasoning extends beyond mere memory lookup. More broadly, meta-reasoning functions as a sense-making process within a knowledge system, enabling people to intuitively direct their learning to discern relations between entities, interact with the same entities in diverse ways, and ground these interactions for both short- and long-term goals. Metacognitive reasoning thus emerges as a foundational mechanism for generalization. In contrast, meta-reasoning systems in artificial intelligence are primarily focused on resource allocation problems. While resource management is essential for efficient functioning, it is crucial to ideate metacognitive generalization, illustrated in Figure 1, to enable generalist behaviors in artificial agents.
@@
Figure 1: a) Illustration of metacognitive reasoning process in a navigation domain, where an agent (depicted as a red triangle) must interpret its environment to achieve the goal of retrieving a blue ball located behind a locked door and various obstacles. This process requires reasoning over abstract object types and their relations, which in turn guide the agent’s interactions with objects. Such abstract reasoning fundamentally relies on higher-order abstract structures and compositional interactions over these abstractions. b) Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities. The metacognitive knowledge-guided approach exploits inherent abstractions and relations between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
Metacognitive reasoning is fundamentally grounded in interaction, requiring agents to draw upon their interpretations of abstract knowledge to guide their interactions. Interpretation precedes and shapes interaction; it involves developing both natural and abstract understandings of the environment that can be deployed as needed to make sense of the current state and plan subsequent interactions. To illustrate this process, consider the example in Figure 1a, which depicts an agent and its interactions in a navigation domain. To build knowledge of the environment, an agent must develop a conceptual understanding of the objects in the environment and how to interact with them. As some objects may even appear similar, the agent must decipher their abstract types. Learning to interact then involves learning which interactions are possible with these abstract object types and the relations between them. Type serves as a natural abstraction, and interpretation thus incorporates this meta-level reasoning over abstract knowledge. The utility of interactions is then learned through meta-level policies in this abstract space, which can subsequently be leveraged to recommend interactions and execute compositions of interactions at a lower level, thereby facilitating problem-solving.
@@
Our proposal is based on three main principles for flexibly solving complex tasks. Natural abstraction: Real-world environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes. Entities in the instance space inherit properties and relations from a higher-level type space. Type-space graphs represent the underlying objects and their relations in environments. This abstraction is where the knowledge resides and it is the knowledge or meta level. Interactions: An agent’s metacognitive process employs the type-space graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals. While the metacognitive policy guides agents based on abstract knowledge, agents interact with entities in the instance space which are instantiations of abstract types. This is the interaction level. Composition of interactions: The agent interacts with meta-guided objects as a recommended interaction type and executes low-level policies corresponding to a guided interaction type. Therefore, a composition of these interactions as interaction pathways for graph traversal at both the meta level and the instance level facilitates execution of low-level actions which are loosely coupled with abstract knowledge through interactions.
@@
Integrating these principles into metacognitive learning, reasoning, and control provides the levers and connecting points for instilling generalist behaviors in agents. Agents’ flexible behaviors stem from their ability to interpret knowledge, learn how to interact with objects through meta level policies and leverage interaction pathways enabling quick adaptation. The proposed metacognitive framework captures the essence of generalist behaviors and may have far-reaching implications across various domains. We describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation in a navigation domain and discuss how we envision it enabling flexibility through meta-level guided interactions.
@@
Compositionality and abstractions are central to generalization learning, with abstraction requiring a mechanism for grounding representations in actual objects. An object-centric model Veerapaneni et al. (2020) was presented in which object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning. Similarly, a symbolic object-centric representation Vijay et al. (2019) was used to study the impact of adding prior knowledge through relations in an object graph. These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making. Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks leveraging natural relational structures. We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is achieved through learning over knowledge graphs with neighborhood aggregation.
@@
Hierarchical reinforcement learning approaches Dayan and Hinton (1992); Sutton et al. (1999); Pateria et al. (2021); Bagaria et al. (2021) learn a higher-level policy by choosing subtasks as higher-level actions, which decompose tasks into subtasks. Higher-level actions persist over a longer timescale, referred to as temporal abstraction Barto and Mahadevan (2003); Dietterich (2000); Sutton et al. (1999). In contrast to options Sutton et al. (1999), which combine a subtask with the core MDP, MAXQ value function decomposition Dietterich (2000) decomposes the core MDP into smaller sub-MDPs, each with a separate learnable policy. Other methods induce learning hierarchies based on teacher signals Portelas et al. (2020); Hutsebaut-Buysse et al. (2022) or a progression from easier to more difficult tasks Campero et al. (2020). These methods enable agents to learn hierarchy of policies, but they remain in the object space for both low- and high-level actions and do not meet a key meta-reasoning demand: type abstractions and loose coupling between abstraction and execution space. Methods such as Ten et al. (2022); Oudeyer et al. (2007) encourage exploration through intrinsic reward or motivation, Klyubin et al. (2005) develop internal goals for empowerment, focusing on aspects different than ours. On the other hand, neuro-symbolic reasoning approaches Hitzler and Sarker (2022); Zhang and Sheng (2024) aim to overcome the limitations of purely symbolic or purely neural methods by integrating them into hybrid models that combine the learning capacity of neural networks with the logical inference and interpretability of symbolic systems. However, significant challenges persist in facilitating interaction and integration between these disparate paradigms to ensure a seamless, flexible, and semantically grounded natural unification.
@@
Our approach meets these demands by leveraging natural higher-order knowledge structures in the form of type space graphs for meta-level policy and facilitating interactions with objects via an intermediary between meta-level and execution space, thus enabling loose coupling and reuse of low-level policies. The appeal of our approach is that interaction policies can be swapped in and out depending on the interaction type. Our work distinguishes itself by incorporating relational abstract knowledge structures to guide interactions that govern low-level execution.
@@
Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of two levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, current meta-reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognitive approaches such as the one proposed here.
@@
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes. Edge from node $u\in\mathcal{V}$ to node $v\in\mathcal{V}$ is given by $(u,v)\in\mathcal{E}$ with attributes $\mathtt{e}_{u,v}$. One typical task is to predict the label or value of the entire graph. Graph neural networks (GNN) (Gilmer et al. (2017); Hamilton et al. (2017) and others) use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$. Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
@@
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which improves static attention problems in graph attention layers.
@@
We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches the state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992), directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Proximal policy optimization (PPO) Schulman et al. (2017) is an on-policy policy optimization technique that addresses concerns in actor-critic methods using clipped objectives. We adapt PPO for learning policies.
Metacognitive framework
@@
Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interpretation of knowledge and interpretation guided interactions with objects. The meta-policy recommends localized agent-object interactions, which the agent executes following interaction policies to construct a compositional solution. Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task-generalization characteristics.
@@
Figure 2: The meta-cognitive generalization framework consists of three levels: knowledge ($\mathcal{K}$), interaction ($\mathcal{I}$), and execution ($\mathcal{X}$), which are based on abstraction and computation for providing generalization capabilities. The knowledge level contains type space knowledge graphs, which represent higher order abstractions. The instance space contains object instance graphs. The interaction level serves as an intermediary, making recommendations for interactions. The execution level is the lowest level at which the agent acts based on the recommended interaction subgoals. Computations proceed through recognition, recommendation, execution, and interactions to collect experiences for updates to interaction and meta policies.
@@
The environment $\Xi$ is partially observable and consists of objects $e=\{e_{1},e_{2},...,e_{n}\}$ and an agent $A$. It is assumed that the agent has a recognition system that can access its partial view $X$ and associate objects with their corresponding types, such that the type of an object $e_{i}\in e$ is in entity types $E=\{E_{1},E_{2},...,E_{m}\}$. The agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ within the environment. When making decisions, the agent organizes its perceived knowledge into graphs: an instance-level graph and a type-level graph. An instance-level graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. A type-level graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their type level or abstract relations $\Lambda$. Relations, whether between objects or entity types, retain the same meaning, such as adjacency indicating that two objects are close to one another. The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and deriving their relationships. The function $\Psi:G_{I}\to G_{K}$ maps instance space graphs to type space graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects. If the agent encounters a novel object without a corresponding type, it must create a new type in its type space or assign the object to a related type. In the framework, $\Psi$ is learnable and represents a belief in mapping objects to types. However, this article treats the mapping as fixed, from object nodes in the instance graph $G_{I}$ to type nodes in the type graph $G_{K}$. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type space graphs.
The agent’s knowledge level is also referred to as the meta-level. At the meta-level, the agent uses a meta-policy that accepts a type space graph and returns interactions or meta-level actions for object interaction. Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$. Similarly, for each interaction type, the agent uses an interaction policy to guide low-level actions $u$ within the environment. Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space graph and available interactions, while the meta or knowledge level $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space graph and meta-level actions or interactions.
@@
After the agent constructs the type and instance graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s type and instance graphs and meta-level policy $\pi$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$. For each object, it adds an activation relation between the agent and the object in both the instance space graph $G_{I}$ and the type space graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction. For $G_{K}$, the agent’s meta-policy returns the state value and the interaction probability. The state value represents the value of activating that specific object. Then, an object $e_{i}$ for interaction is proposed by sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object. Therefore, at each step, the meta-level of the agent activates a candidate object $e_{i}$ for interaction $\lambda_{p}$, setting the agent’s interaction goal $g^{i}_{p}=(e_{i},\lambda_{p})$. Acting on the recommended interaction goal, the agent begins interacting with the environment. After each round of interaction, the interaction status is passed to the meta-level, which then proposes the next candidate object and interaction. This process continues until the task is completed or the environment is terminated.
@@
Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge-type encoding for different types of relations. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,

### v4

# Knowledge-Centric Metacognitive Learning

@@ Abstract
Interactions are central to intelligent reasoning and learning abilities, with the interpretation of abstract knowledge guiding meaningful interaction with objects in the environment. While humans readily adapt to novel situations by leveraging abstract knowledge acquired over time, artificial intelligence systems lack principled mechanisms for incorporating abstract knowledge into learning, leading to fundamental challenges in the emergence of intelligent and adaptive behavior. To address this gap, we introduce knowledge-centric metacognitive learning based on three key principles: natural abstractions, knowledge-guided interactions through interpretation, and the composition of interactions for problem solving. Knowledge learning facilitates the acquisition of abstract knowledge and the association of interactions with knowledge, while object interactions guided by abstract knowledge enable the learning of transferable interaction concepts, abstract reasoning, and generalization. This metacognitive mechanism provides a principled approach for integrating knowledge into reinforcement learning and offers a promising pathway toward intelligent and adaptive behavior in artificial intelligence, robotics, and autonomous systems.
@@ Introduction
A hallmark of human intelligence is the capacity to construct and interpret abstract knowledge. Interpretation grounded in such abstractions enables meaningful interaction with the environment, forming the foundation for knowledge learning, flexible planning, and decision-making. It is fascinating how effortlessly people discern natural abstractions and leverage them to interact with their environments in flexible and adaptive ways. In contrast, contemporary artificial intelligence systems lack the principled integration of abstract knowledge into learning, resulting in fundamental limitations in the emergence of intelligent behavior. For example, autonomous systems trained solely on object-level data often fail to respond to changes in world characteristics due to the absence of integrated knowledge learning. Such gaps in knowledge learning lead to non-robust, unreliable, and specialized behavior in unknown and complex scenarios. Machine learning methods Goodfellow et al. (2016); Bishop and Nasrabadi (2006) learn a mapping function by optimizing an objective between predictions and targets; however, these targets lack the underlying know-how required for world understanding. In reinforcement learning Sutton and Barto (2018), agents collect experiences through interaction with their environments and learn policies for action selection, yet even model-based approaches, despite learning a dynamics model, fail to integrate abstract knowledge into learning and world models. Learning, at its core, is knowledge learning for an abstract understanding of the world and responding adaptively through the use of such knowledge. Addressing these gaps necessitates a knowledge-centric metacognitive learning approach.
Meta-reasoning about subsequent actions underpins flexible decision-making. More broadly, meta-reasoning acts as a sense-making process within knowledge systems, enabling people to intuitively direct their learning to discern abstract relational structure and ground object interactions. However, meta-reasoning in artificial systems predominantly acts as a resource controller, monitoring lower-level resources and managing resource allocation for efficient functioning. While such resource management is essential, a resource-controller–only view is severely limiting for facilitating abstract reasoning and adaptive behaviors. Meta-reasoning must pivot toward interpretation-based reasoning within abstract knowledge. This paradigm of abstract interpretive reasoning and learning – termed metacognitive reasoning and learning – is central to much of knowledge learning, interaction with objects in the physical world, and manifested in the emergence of autonomous intelligent behaviors. Figure 1 illustrates this process in a navigation domain. To develop an understanding of the world, an agent must learn abstract knowledge of objects where object types serve as natural abstractions, their relations, and the ways in which they can be interacted with. Interpretation over this abstract knowledge guides interactions enabling flexible problem solving.
@@
Figure 1: a) Illustration of metacognitive approach in a navigation domain, where an agent must develop an abstract understanding of its environment and interpret it to achieve the goal of retrieving a blue ball located behind a locked door, blocked by obstacles, with the key hidden inside a box. b) A knowledge centric metacognitive learning approach exploits inherent object abstractions and relations to generate interaction goals and compose interactions, thereby enabling the emergence of intelligent and flexible behavior.
@@
Our proposal is based on three main principles for knowledge-centric learning and problem-solving. Natural abstraction: Real-world environments are naturally composed of entities categorized by their types, similar to an inherent hierarchy of objects and classes. Objects in the instance space inherit properties and relations from a higher-level type space. Type space graphs provide an abstract representation of objects and their relationships within the environment. The abstraction space is where the knowledge resides and is referred to as the knowledge or meta level. Interactions: An agent’s metacognitive process employs type space graphs to create plans that specify which entities to interact with and the nature of these interactions in order to achieve its goals. While the meta-level policy guides the agent based on abstract knowledge, the agent executes interactions in the instance space, where objects are type instantiations. This constitutes the interaction level. Composition of interactions: The meta-level guides the agent to interact with recommended objects using interaction types, and the agent executes low-level policies corresponding to guided interactions. The composition of these interactions forms interaction pathways for graph traversal at both the meta and instance level, thereby facilitating the execution of low-level actions. The process is loosely coupled through interactions, enabling flexible and adaptive behavior.
@@
Integrating these principles into metacognitive learning provides foundational mechanisms for endowing agents with knowledge learning and adaptive capabilities, arising from their ability to interpret abstract knowledge, learn object interactions guided by meta-level policies, and leverage interaction pathways for adaptation. The proposed metacognitive learning approach captures the essence of abstract reasoning, knowledge learning, and flexible behavior, with potentially far-reaching implications across domains that require abstract reasoning and interaction with the physical world. In the remaining of this article, we describe the key components and exemplifying characteristics of the approach, followed by an experimental evaluation in a navigation domain, and discuss how it enables adaptive problem solving through knowledge–guided interaction learning.
@@
Compositionality and abstractions are central to generalization, with abstraction requiring a mechanism for grounding representations in objects. An object-centric model Veerapaneni et al. (2020) was presented in which object representations were constructed from raw observations and local functions with uniformity were then used for prediction and planning. A symbolic object-centric representation Vijay et al. (2019) was used to study the impact of adding prior knowledge through relations in an object graph. These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making. Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks for problem solving. We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is achieved through learning over knowledge graphs with neighborhood aggregation.
@@
Hierarchical reinforcement learning approaches Dayan and Hinton (1992); Sutton et al. (1999); Pateria et al. (2021); Bagaria et al. (2021) learn a higher-level policy by choosing subtasks as higher-level actions, which decompose tasks into subtasks. Higher-level actions persist over a longer timescale, referred to as temporal abstraction Barto and Mahadevan (2003); Dietterich (2000); Sutton et al. (1999). In contrast to options Sutton et al. (1999), which combine a subtask with the core MDP, MAXQ value function decomposition Dietterich (2000) decomposes the core MDP into smaller sub-MDPs, each with a separate learnable policy. Other methods induce learning hierarchies based on teacher signals Portelas et al. (2020); Hutsebaut-Buysse et al. (2022) or a progression from easier to more difficult tasks Campero et al. (2020). These methods enable agents to learn hierarchy of policies, but they remain in the object space for both low- and high-level actions and do not meet key meta-reasoning demands: type abstractions and loose coupling. Methods such as Ten et al. (2022); Oudeyer et al. (2007) encourage exploration through intrinsic reward or motivation, Klyubin et al. (2005) develop internal goals for empowerment, focusing on aspects different than ours.
Goal-conditioned reinforcement learning Liu et al. (2022); Pong et al. (2018) seeks to support multi-task generalization by conditioning policies on goals, yet remains challenged due to the lack of integrated knowledge based reasoning and learning in complex tasks. We address these limitations through a metacognitive approach, in which abstract knowledge constrains learning, generates and guides interactions, enabling flexible behavior. Neuro-symbolic approaches Hitzler and Sarker (2022); Zhang and Sheng (2024) aim to overcome the limitations of purely symbolic or neural methods by integrating them into hybrid models that combine the learning capacity of neural networks with the logical inference and interpretability of symbolic systems. However, significant challenges persist in facilitating interaction and integration between these disparate paradigms to ensure a seamless and semantically grounded unification.
@@
Our approach addresses these demands by leveraging natural higher-order knowledge structures, represented as type-space graphs, for meta-level policy learning, and by mediating object interactions via an intermediary between the meta and execution levels, thus enabling loose coupling and reuse of low-level policies and interaction policies can be swapped in and out based on interaction type. Our work is distinguished by its principled incorporation of relational abstract knowledge to guide interactions that govern low-level execution.
@@
Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, current meta-reasoning, primarily as a resource manager, is insufficient for exhibiting intelligent and generalist behaviors, necessitating a meta-cognitive approach.
@@
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes. Edge from node $u\in\mathcal{V}$ to node $v\in\mathcal{V}$ is given by $(u,v)\in\mathcal{E}$ with attributes $\mathtt{e}_{u,v}$. One typical task is to predict the label or value of the entire graph. Graph neural networks (GNN), Gilmer et al. (2017); Hamilton et al. (2017) and others, use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$. Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
@@
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer k1k-1, $\bigoplus$ is a differentiable permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which improves static attention problems in graph attention layers.
@@
We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches the state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992), directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity A(s,a)=Q(s,a)V(s)A(s,a)=Q(s,a)-V(s) can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Policies are learned by adapting Proximal policy optimization (PPO) Schulman et al. (2017), an on-policy policy optimization that addresses concerns in actor-critic methods using clipped objectives.
Metacognitive Learning Framework
@@
Our proposal for metacognitive learning centers on the interpretation of abstract knowledge to guide interactions with objects. A meta-policy recommends localized agent–object interactions, which the agent executes via interaction policies to construct a compositional task solution. Here, we outline the main ideas of the metacognitive learning framework, Knowledge-Interaction-eXecution (KIX) (see Figure 2).
@@
Figure 2: The metacognitive learning framework comprises three levels: knowledge or meta ($\mathcal{K}$), interaction ($\mathcal{I}$), and execution ($\mathcal{X}$), based on abstraction and computation. The knowledge level contains type space knowledge graphs that represent higher-order abstractions, while the instance space comprises object instance graphs. The interaction level serves as an intermediary, making recommendations for interactions that the agent executes at the execution level. Computation proceeds through cycles of recognition, recommendation, execution, and interaction, to collect experiences for updates to both the meta-level and interaction policies.
@@
The environment $\Xi$ is partially observable and consists of objects $e=\{e_{1},e_{2},...,e_{n}\}$ and an agent $A$. It is assumed that the agent has a recognition system that can access its partial view $X$ and associate objects with their corresponding types, such that the type of an object $e_{i}\in e$ is in entity types $E=\{E_{1},E_{2},...,E_{m}\}$. The agent can perform low-level actions $u=\{u_{1},u_{2},$...\} within the environment. When making decisions, the agent organizes its knowledge into instance, and type graphs. An instance graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. A type graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their abstract relations $\Lambda$. Relations, whether between objects or entity types, retain the same meaning, such as adjacency indicating that two objects are close to one another. The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and relations. The function $\Psi:G_{I}\to G_{K}$ maps instance graphs to type graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects. If the agent encounters a novel object, it must create a new type in its type space or assign the object to a related type. In the framework, $\Psi$ is learnable and represents a belief, mapping objects to types. However, this article treats the mapping as fixed, from object nodes in the instance graph $G_{I}$ to type nodes in the type graph $G_{K}$. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type graphs.
The agent’s knowledge level is also referred to as the meta-level. At the meta-level, the agent uses a meta-policy that accepts a type graph and returns interactions or meta-level actions for object interaction. Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$. For each interaction type, the agent uses an interaction policy to execute low-level actions $u$ within the environment. Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises its instance graph and available interactions, while the knowledge or meta level $\mathcal{K}=(A,G_{K},\lambda)$ comprises its type graph and meta level actions or interactions.
@@
After the agent constructs the type and instance graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s type and instance graphs and meta-level policy $\pi$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$. For each object, it adds an activation relation between the agent and the object in both the instance graph $G_{I}$ and the type graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction. For $G_{K}$, the agent’s meta-policy returns the state value and the interaction probability. The state value represents the value of activating that specific object. Then, an object $e_{i}$ for interaction is proposed by sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object. Therefore, at each step, the meta-level of the agent activates a candidate object $e_{i}$ for interaction $\lambda_{p}$, setting the agent’s interaction goal $g^{i}_{p}=(e_{i},\lambda_{p})$. Acting on the recommended interaction goal, the agent begins interacting with the environment. After each round of interaction, the interaction status is passed to the meta-level, which then proposes the next candidate object and interaction. This process continues until the task is completed or the environment is terminated.
@@
Meta-level representations of type graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge-type encoding for different types of relations. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,